Calibration Runs

Calibration Runs


  Run 5063103 Pedestal

  Run 5063104   2 MOhm in IFC east (normal)

  Run 5063105   2 Mohm in IFC west

  Run 5063106   2 Mohm back in IFC east


That was email from Blair on March 4, 2004.


_________________________

There was also:

January 22, 2004:


 Magnet on, B polarity

 No beam in RHIC

 Local clock

 All laser beams

 Gated grid was varied on both east and west


  Run 5022005   Pedestal

  Run 5022006   GG = - 115 (Nominal)

  Run 5022007   GG = -140

  Run 5022008   GG = -165

  Run 5022009   GG = - 90

  Run 5022010   GG = -65


   For the last run 4 outer, 8 outer, 22 outer, 24 outer

    and 14,15 inner had a yellow alarm

   I looked at each and they were at ~ 63 volts. According

to Geno, that difference shouldn't be noticeable.


________________

Set 2 was taken at ~20040122.164000 GMT

DB has altered GG values for these.


Set 1 was taken at ~20040303.213000 GMT

DB does not have altered field cage short values for these.